OpenStudy (anonymous):

can you help me figure out how 4x^2 - 4y^2 + 8y - 4x - 11 = 0 is an ellipse? Steps will be appreciated.

OpenStudy (anonymous):

4x^2-4x-4y^2+8y=11 (2x-1)^2-1-4(y^2-2y)=11 (2x-1)^2-4((y-1)^2-1)=12 (2x+1)^2-4(y-1)^2+4=12 (2x+1)^2-4(y-1)^2=8